Find the vertex, line of symmetry, the maximum or minimum value of the quadratic equation, and graph the function. f(x)= -2x^2+2x+6: Find the vertex, line of symmetry, the maximum or minimum value of the quadratic equation, and graph the function. f(x)= -2x^2+2x+6
Answer by Alan3354(1934)

Find the vertex, line of symmetry, the maximum or minimum value of the quadratic equation, and graph the function. f(x)= -2x^2+2x+6
-----------
set f(x) = 0 to find where it crosses the x-axis.
-2x^2+2x+6=0
x^2 - x - 3 = 0
Solved by pluggable solver: SOLVE quadratic equation (work shown, graph etc)
Quadratic equation ax^2+bx+c=0 (in our case 1x^2+-1x+-3 = 0) has the following solutons:

x[12] = (b+-sqrt( b^2-4ac ))/2\a

For these solutions to exist, the discriminant b^2-4ac should not be a negative number.

First, we need to compute the discriminant b^2-4ac: b^2-4ac=(-1)^2-4*1*-3=13.

Discriminant d=13 is greater than zero. That means that there are two solutions:  x[12] = (--1+-sqrt( 13 ))/2\a.

x[1] = (-(-1)+sqrt( 13 ))/2\1 = 2.30277563773199
x[2] = (-(-1)-sqrt( 13 ))/2\1 = -1.30277563773199

Quadratic expression 1x^2+-1x+-3 can be factored:
1x^2+-1x+-3 = (x-2.30277563773199)*(x--1.30277563773199)
Again, the answer is: 2.30277563773199, -1.30277563773199. Here's your graph:
graph( 500, 500, -10, 10, -20, 20, 1*x^2+-1*x+-3 )

The vertex is at the minimum: set the 1st derivative to 0
2x-1=0
x = 1/2
---------
sub for x in the eqn
y = (1/4) - 1/2 - 3
y = -3 1/4 = -13/4
So the vertex is at (1/2,-13/4)
Since there's no xy term, the axis of symmetry is parallel to the y-axis. It goes thru the vertex, so it's:
x = 1/2